# Constraining the Cosmic-ray Energy Based on Observations of Nearby Galaxy Clusters by LHAASO

### v1

@@ Appendix B EBL model
In this appendix, we provide detailed information regarding the EBL model. $\gamma$-rays within the energy range observed by LHAASO undergo attenuation due to pair-production interactions with the CMB and EBL as they traverse extragalactic space. The photon-photon pair-production cross-section averaged over directions of the background-radiation field depends on the product of energies of colliding photons. For the given energy of the $\gamma$-ray photon $E_{\gamma}$, it peaks at the wavelength of background photons $\lambda\sim 1(E_{\gamma}/{\rm 1\ TeV})^{-1}\mu\rm m$. Thus, the $\gamma$-ray opacity above and below 100 TeV is due to the absorption on CMB and EBL, respectively. The opacity has been calculated by performing the line-of-sight integral of the product of the pair production cross section with the energy density of the radiation fields. The dependence of the pair production cross section on the energy is given in Gould & Schreder (1967). The EBL energy density is taken from the Franceschini et al. (2008). Figure B.1 illustrates the variation in absorption efficiency with energy at the respective distances of the Coma, Perseus, and Virgo clusters, as predicted by the EBL model adopted in this paper.

### v2

@@ Appendix B EBL model
In this appendix, we provide detailed information regarding the EBL model. $\gamma$-rays within the energy range observed by LHAASO undergo attenuation due to pair-production interactions with the CMB and EBL as they traverse extragalactic space. The photon-photon pair-production cross-section averaged over directions of the background-radiation field depends on the product of energies of colliding photons. For the given energy of the $\gamma$-ray photon $E_{\gamma}$, it peaks at the wavelength of background photons $\lambda\sim 1(E_{\gamma}/{\rm 1\ TeV})^{-1}\mu\rm m$. Thus, the $\gamma$-ray opacity above and below 100 TeV is due to the absorption on CMB and EBL, respectively. The opacity has been calculated by performing the line-of-sight integral of the product of the pair production cross section with the energy density of the radiation fields. The dependence of the pair production cross section on the energy is given in gouldPairProductionPhotonPhoton1967a. The EBL energy density is taken from the franceschiniExtragalacticOpticalinfraredBackground2008. Figure B.1 illustrates the variation in absorption efficiency with energy at the respective distances of the Coma, Perseus, and Virgo clusters, as predicted by the EBL model adopted in this paper.